OpenStudy (anonymous):

Please help! A4 = {1, (123), (132), (124), (142), (134), (143), (234), (243), (12)(34), (13)(24), (14)(23)} is the alternating group, which consists of all even permutations of S4. Consider the following subset H1 = {1, (12)(34), (13)(24), (14)(23)} of S4. Is H1 a subgroup of A4? The same question for the subset H2 = {1,(123), (132), (124), (142), (134), (143), (234), (243)}

OpenStudy (jamesj):

So two questions: 1. Is H1 a subgroup at all? 2. Is it a subgroup of A4? First H1 is closed under multiplication as every element is of order 1 or 2. Second, every member of H1 is also a member of A4 so yes. For H2, it is a subgroup, because the elements are all of order 1 or 3, and written pairwise next to their inverse e.g., (123)^2 = (123)^-1 = (132) Now is it subgroup of A4?

OpenStudy (anonymous):

Ok, I'm a little confused. One of the axioms for the definition I have says "for all a in H, a^-1 is in H. So doesn't this mean that H1 is not a subgroup since we can't find all the inverses of all the elements (belonging to H1) in H1?

OpenStudy (jamesj):

What's the inverse of (12)(34)?

OpenStudy (anonymous):

(21)(43)?

OpenStudy (jamesj):

and (21)(43) = (12)(34)

OpenStudy (jamesj):

yes?

OpenStudy (anonymous):

yes.

OpenStudy (jamesj):

So every element in H1 is its own inverse. e of course is of order 1, e^1 = e; all the others are of order 2: e.g., ( (12)(34) )^2 = e

OpenStudy (jamesj):

Hence for every h in H1, h^-1 is also in H1.

OpenStudy (jamesj):

...because h^-1 = h

OpenStudy (anonymous):

Ok. So is H2 also a subgroup of A4?

OpenStudy (jamesj):

Go through the axioms 1. Is it a (sub)group: identity, closure under multiplication, existence of inverses 2. Then is it a subset of A4?

OpenStudy (jamesj):

Are the answers yes to all those questions?

OpenStudy (anonymous):

Yes. i think? Can you explain why you use the order of the cycles to figure out whether it's closed under multiplication?

OpenStudy (jamesj):

it's a short-hand really. We don't need to use it, but it's useful to know and after working with these things for enough time, you'll build up more intuition about them and think about them in this among other ways.

OpenStudy (anonymous):

ok. Thank you.

OpenStudy (anonymous):

Wait, was 'yes' the correct answer?

OpenStudy (jamesj):

Yes, H2 is also a subgroup of A4 as you've written it. Now my question for you: is that really A4? Why is (132) for instance an even permutation?

OpenStudy (anonymous):

Honestly, I don't know the answer to that question. The problem was just written that way. I think a permutation is even if the 'sign' is even?

OpenStudy (jamesj):

It is A4. Anyway, let's see what your teacher/professor gives you next week. I think s/he is setting you up for something interesting about A4.

OpenStudy (jamesj):

ok ... I need to go do my own stuff now. Good luck with group theory. It's nice stuff; I hope you're enjoying it.

OpenStudy (anonymous):

It is pretty interesting. Thanks so much for your help!